OpenStudy (anonymous):

F(s)=5s^2+34s+53/(s+3)^2(s+1) determine inverse laplace transform will format in question!

OpenStudy (anonymous):

\[F(s)=5s^2+34s+53/(s+3)^2(s+1)\]

OpenStudy (anonymous):

I did... \[A/(s+3)+Bs+C/(s+3)^2+D/(s+1)\]

OpenStudy (anonymous):

Just spread out 53/(s+3)2(s+1) and then use the inverse rules to do the rest. This one is pretty straight forward.

OpenStudy (anonymous):

but then got stuck at \[A(s+3)^2(s+1)+Bs+C(s+3)(s+1)+D(s+3)^2(s+3)\]

OpenStudy (anonymous):

am I doing this right?

OpenStudy (anonymous):

if so I just need to learn binomial expansion cause im pretty sure that's all this step is

OpenStudy (anonymous):

@oldrin.bataku any advice?

OpenStudy (anonymous):

just a simple yes im on the right track will suffice then ill watch youtube videos on binomial expansion to try and continue

OpenStudy (anonymous):

you multiplied to get rid of the denominators incorrectly :-/

OpenStudy (anonymous):

also why the \(Bs\) term?

OpenStudy (psymon):

You would only need that Bx + C numerator if the factor itself is quadratic like (s^2 +3). The squared term on the outside doesn't make it Bs + C, it'd still just be B for that factor.

OpenStudy (anonymous):

oh ok thank you @Psymon I will try it again

OpenStudy (abb0t):

Are you using partial fraction decomposition?

OpenStudy (anonymous):

@abb0t ya

OpenStudy (abb0t):

I don't think you need \(\frac{Bs+C}{(s-3)^2}\) I think. Can't you just use \(\frac{B}{(s-3)^2}\)?

OpenStudy (psymon):

Yeah, it should just be B in the numerator.

OpenStudy (abb0t):

Oh nvm. I just read that. My bad. Too dam early for ODE.

OpenStudy (psymon):

Agreed, haha.

OpenStudy (anonymous):

the reason that works is because we just split into \((s+3),(s+3)^2\) terms individually :-p observe:$$\frac{A}{s+3}+\frac{B}{(s+3)^2}=\frac{A(s+3)}{(s+3)^2}+\frac{B}{(s+3)^2}=\frac{As+(3A+B)}{(s+3)^2}$$

OpenStudy (anonymous):

ok I got A=19/6 B=34/3 and C=-19/6 how do I determine what is what from \[L^{-1}[(19/6)/(S+3)+(34/3)/(S+3)^2-(19/6)/(S+1)]\] the solution is \[-e^{-3t}+2te^{-3t}+6e^{-t}\] so I think my a,b,and c are right but not sure.... @oldrin.bataku

OpenStudy (anonymous):

$$F(s)=\frac{5s^2+34s+53}{(s+3)^2(s+1)}=\frac{A}{s+3}+\frac{B}{(s+3)^2}+\frac{C}{s+1}\\A(s+3)(s+1)+B(s+1)+C(s+3)^2=5s^2+34s+53$$Let \(s=-1\) and our equation reduces to:$$C(-1+3)^2=5(-1)^2+34(-1)+53\\4C=5-34+53\\4C=24\\C=6$$Let \(s=-3\) and we get:$$B(-3+1)=5(-3)^2+34(-3)+53\\-2B=45-102+53\\2B=4\\B=2$$To find our last coefficient consider:$$B(s+1)+C(s+3)^2=2(s+1)+6(s+3)^2=2s+2+6(s^2+6s+9)=6s^2+38s+56$$and also:$$(s+3)(s+1)=s^2+4s+3$$hence it makes intuitive sense that \(A=-1\) yielding in conclusion:$$F(s)=-\frac1{s+3}+\frac2{(s+3)^2}+\frac6{s+1}$$

OpenStudy (anonymous):

now, we take the inverse transform to return to time domain:$$f(t)=\mathcal{L}^{-1}\left\{-\frac1{s+3}+\frac2{(s+3)^2}+\frac6{s+1}\right\}=-e^{-3t}+2te^{-3t}+6e^{-t}$$

OpenStudy (anonymous):

@oldrin.bataku thank you so much im sure that this will help me finish this sections homework. just wondering what your major is and where you go to school???

OpenStudy (anonymous):

in high school but I hope to study engineering and math

OpenStudy (anonymous):

@SithsAndGiggles can you explain to me how he got A cause apparently it isn't coming to me intuitively :(

OpenStudy (anonymous):

I get that (s+3)(s+1) comes from A and the other comes from b and c with the numbers we found....but that's it

OpenStudy (anonymous):

i tried adding them together and set equal to the numerator but that didn't help me

OpenStudy (anonymous):

\[A(s+3)(s+1)+B(s+1)+C(s+3)^2=5s^2+34s+53\] I'll solve for the constants using a different method (I usually make mistakes when I use oldrin's method) First thing to do would be to expand all the binomial expressions: \[As^2+4As+3A+Bs+B+Cs^2+6Cs+9C=5s^2+34s+53\] Group the terms with common powers of \(s\): \[(A+C)s^2+(4A+B+6C)s+(3A+B+9C)=5s^2+34s+53\] Match up terms containing same powers of \(s\) on the other side of the equation. Doing so yields the system \[\begin{cases}A+C=5\\4A+B+6C=34\\3A+B+9C=53\end{cases}\] Solving yields \(A=-1,~B=2,~C=6\). I'll leave the details of solving the system to you... Anyway, I get the same results as oldrin above.

OpenStudy (anonymous):

thanks for helping!

OpenStudy (anonymous):

I just considered that since I had \(6s^2\) without the \(A\) term and wanted \(5s^2\) that we'd need \(A=-1\) to since \(6+-1=5\)